OpenStudy (anonymous):

CHECK MY WORK! (If you'd like) What is the sum of an 8-term geometric series if the first term is 13 and the last term is -3,639,168? Answer: -3,119,285 What is the sum of a 7-term geometric series if the first term is -11, the last term is -171,875, and the common ratio is -5? Answer: -36,047 What is the 6th term of the geometric sequence where a1 = 1,024 and a4 = -16? Answer: -0.25

OpenStudy (anonymous):

@satellite73 help ;-; this time im confused :P

OpenStudy (campbell_st):

well you need to know \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] so Question 1 seems to be missing something, perhaps the common ratio... for question 2 you should have \[S_{7} = \frac{-11 \times (1 - (-5)^7)}{1 - (-5)}\] which gives a different answer than the 1 you posted

OpenStudy (anonymous):

for the first one \(r=\sqrt[7]{\frac{-3,639,168}{13}}=-6\)

OpenStudy (campbell_st):

the last question \[-16 = 1024 \times r^{4 - 1}\] you need to find r 1st then find the 6th term

OpenStudy (anonymous):

Wow thank you guys! Ok, so for the second question I got 144,177.

OpenStudy (anonymous):

i get \(3,119,285\) for the first one same answer you got but mine is positive

OpenStudy (anonymous):

getting \(r=-6\) the sum should be \[\frac{13(6^8-1)}{-6-1}\] or oops i think you are right and i am wrong!

OpenStudy (anonymous):

your answer is correct, my answer was off by a minus sign

OpenStudy (campbell_st):

I got 143231 so you better check it

OpenStudy (anonymous):

@campbell_st i mean for the first one

OpenStudy (anonymous):

@campbell_st I got 143,231 but my multiple choice has only 144,177. So I assumed that it was the answer.

OpenStudy (campbell_st):

lol... sorry yes... it should be negative... and the 2nd should also be negative

OpenStudy (anonymous):

No, no, no don't be sorry! Seriously you two are amazing! :)

OpenStudy (campbell_st):

ummm don't know what to say about the 2nd question...

OpenStudy (anonymous):

Ok so the final answers are: Question 1: -3,119,285 Question 2: 144,177 Question 3: -1

OpenStudy (anonymous):

yeah the 2nd question is weird. I've double-checked to see if I typed it in wrong and I haven't. I'll let you know what the answer to that question is when I submit this test. :)

OpenStudy (campbell_st):

oops last question \[a_{6} = 1024 \times (\frac{-1}{4})^5\] so your answer of -1 is correct

OpenStudy (anonymous):

awesome! :) Thank you so much!